How do you solve by substitution or elimination these linear equations?
{2x+y=5}
{1x+2y=4}
I get confused on these steps.

Answer by mananth(16946)

2x+1y=5 .............1
1x+2y = 4 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-4x-2y= -10
1x+2y= 4
Add the two equations
-3x=-6
/-3
x=2
plug value of x in (1)
2x+1y=5
4+1y=5
1 y = 5 -4
y = 1